OpenStudy (anonymous):

plz help fan medal and testimonial which of the following are continuus for all real values of x? I. f(x)= x^2+5/x^2-1 ii. g(X)= 3/x^2+1 iii. h(X)={x-1} choices- ii and iii only I and ii only I only ii only

OpenStudy (anonymous):

I know for a fact that choice b is wrong.

OpenStudy (anonymous):

@dan815 @Math2400 @Camila1315 @notyourdroid

myininaya (myininaya):

what does { } mean?

myininaya (myininaya):

for the two fractions you can find where they are not continuous by finding when the denominator is zero

OpenStudy (anonymous):

parenthesis

myininaya (myininaya):

so h(x)=x-1?

OpenStudy (anonymous):

yes

myininaya (myininaya):

h(x)=x-1 is just a line then

myininaya (myininaya):

do you have any questions?

OpenStudy (anonymous):

so my suspicion was right lol h(X)=x-1 is my answer? lol

myininaya (myininaya):

I didn't say that

myininaya (myininaya):

did you find when the bottoms were zero (if any values) for the first two?

OpenStudy (anonymous):

there continuous but I put I and ii as my answer before and I was told it was wrong.

myininaya (myininaya):

I assume the first fraction is: \[f(x)=\frac{x^2+5}{x^2-1} \\ \text{ can you solve } x^2-1=0 \\ \text{ the second fraction is } g(x)=\frac{3}{x^2+1} \\ \text{ can you solve } x^2+1=0\]

OpenStudy (anonymous):

x=1 for both

myininaya (myininaya):

x^2-1=0 x^2=1 x=1 or x=-1 right? but how is 1^2+1 zero?

myininaya (myininaya):

so f is discontinuous at x=1 and also x=-1 but what do you notice about x^2+1=0?

OpenStudy (anonymous):

same as the first.

myininaya (myininaya):

no (1)^2+1 is not zero (-1)^2+1 is not zero

myininaya (myininaya):

both (1)^2+1 and (-1)^2+1 is 1+1 which is 2

myininaya (myininaya):

if I write x^2+1=0 as x^2=-1 If i try to think of a positive number to plug in...I'm gonna have to square it but you know a positive times a positive =? or if i try to think of a negative number to plug in...I'm gonna have to square it but you know a negative times a negative=? or if I plug in 0 which still isn't gonna work because 0^2 is not -1

OpenStudy (anonymous):

okay...

myininaya (myininaya):

I'm asking you to tell me what a positive times a positive will give and what a negative times a negative will give

ganeshie8 (ganeshie8):

{x} usually refers to the fractional part of x

OpenStudy (anonymous):

+ x + = + - x - = +

myininaya (myininaya):

right x^2 is always positive or 0 -1 is negative there is no way x^2 can even be -1 over the real numbers

myininaya (myininaya):

there is no real solution to x^2+1=0

myininaya (myininaya):

therefore g=3/(x^2+1) has no discontinuities ok so h(x)=x-1 or h(x)={x-1} @magy33 ?

myininaya (myininaya):

I'm asking since @ganeshie8 pointed out { } means fractional part

OpenStudy (anonymous):

h(x)= x-1 but does the {} really have anything to do with the equation itself? I askto be sure for next time :)

myininaya (myininaya):

If it is there, it does

ganeshie8 (ganeshie8):

http://mathworld.wolfram.com/FractionalPart.html

OpenStudy (anonymous):

ok. I think then that from my understanding the answer is ii and iii

ganeshie8 (ganeshie8):

I'm inclining toward ii only

myininaya (myininaya):

me to if h={x-1} and not x-1

ganeshie8 (ganeshie8):

the fractional function `h(x) = {x-1}` is discontinuous at every integer

OpenStudy (anonymous):

lol I was half right. thnks for the help guys. lol I still have more questions to go before im done lol I see if I can answer if not ill post lol